OpenStudy (anonymous):

write the general form of the equation of the circle. center: (0,0) radius: 3 I got: sqrt(x^2 + y^2) = 3

OpenStudy (anonymous):

x^2 + y^2 = 9

OpenStudy (anonymous):

Nomss, you are correct, as long as you append: \[\pm \sqrt{x^2+ y^2} = 3\]

OpenStudy (anonymous):

\[\sqrt{x^2+y^2}=3\] is only the upper half of the circle

OpenStudy (anonymous):

the standard form is what i wrote

OpenStudy (anonymous):

him1618 is exactly right.

OpenStudy (anonymous):

their answers are correct :) good luck

OpenStudy (anonymous):

actually this is what i had too: x^2 + y^2 = 9 thanks guys :)

OpenStudy (anonymous):

um no no. i need general form...

OpenStudy (anonymous):

generel form ? can u explain more ?

OpenStudy (anonymous):

Ax^2 + Ay^2 + Dx + Ey + F = 0 A != 0

OpenStudy (anonymous):

so you got it by yourself !

OpenStudy (anonymous):

Yeah, \[x^2 + y^2 = 9\] is the general form

OpenStudy (anonymous):

Good job bro =D !